OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My equation is f(x)=2(x-9)^2+17

OpenStudy (anonymous):

so the average rate of change between two points of a function is the slope between the two points. Remember slope is:\[\frac{ \Delta y }{ \Delta x }\] (change in y over change in x) which can be written as:\[\frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] in this case:\[x_{2}=7\] and \[x _{1}=4\] To find the y-coordinates plug in the x values into f(x) so:\[y _{2}=2(7-9)^2+1\] \[y _{1}=2(4-9)^2+1\] Now just simplify those and plug them into the slope formula to find the average rate of change.

OpenStudy (anonymous):

oh actualy those ones are 17s right?

OpenStudy (anonymous):

sorry replace those +1s with +17s

OpenStudy (anonymous):

im so confused my euqation was f(x)=2(x-7)^2+17